OpenStudy (anonymous):

Evaluate P(3, 0). 0 1 6

OpenStudy (anonymous):

6 Right?

OpenStudy (zarkon):

are you doing permutations?

OpenStudy (anonymous):

yes sirr

OpenStudy (zarkon):

\[P(3,0)=_3P_0=\frac{3!}{(3-0)!}\]

OpenStudy (zarkon):

\[\frac{3!}{3!}=6/6=1\]

OpenStudy (anonymous):

thank u

OpenStudy (solomonzelman):

P(3,0) doesn't really make sense though. You are choosing zero things out of (some) 3 things.

OpenStudy (solomonzelman):

Well, it is a bit silly